# Quantised ${\mathfrak{sl}}_{2}$-differential algebras

### v1

@@ Abstract.
A. Krutov was supported by the GAČR projects 24-10887S and by HORIZON-MSCA-2022-SE-01-01 CaLIGOLA. P. Pandžić was supported by the project “Implementation of cutting-edge research and its application as part of the Scientific Center of Excellence for Quantum and Complex Systems, and Representations of Lie Algebras”, PK.1.1.02, European Union, European Regional Development Fund. This article is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology, www.cost.eu).
@@ 1. Introduction
Let ${\mathfrak{g}}$ be a Lie algebra. H. Cartan introduced the notion of ${\mathfrak{g}}$-differential algebras as a generalisation of differential forms on manifolds with ${\mathfrak{g}}$-action, [Car1, Car2]. Later ${\mathfrak{g}}$-differential algebras appeared in the study of equivariant cohomology [GS, AM1], in Chern–Weil theory [AM2, Mei], and in relation to (algebraic) Dirac operators and Vogan’s conjecture [AM1, HP1, HP2].
There have been several attempts to generalise the notion of ${\mathfrak{g}}$-differential algebras to the setting of quantum groups and noncommutative geometry; for example, see [AC, SWZ, AS]. These works, however, assumed that we either work with a triangular Hopf algebra, or start with a bicovariant calculus on a quantum group, so they do not directly apply the setting of $U_{q}({\mathfrak{sl}}_{2})$ since it is only a quasitriangular, see [Dri1, Dri2], and its bicovariant differential calculus does not have classical dimension [Wor], see also [Jur] for the general case.
In this paper we propose a definition of quantised ${\mathfrak{sl}}_{2}$-differential algebras and give first examples, certain quantised Clifford and exterior algebras. The advantage of our approach is that we start with the quantum exterior algebra defined by Berenstein and Zwicknagl [BZ] of the classical dimension instead of a bicovariant calculus. We use the coboundary structure on the category of $U_{q}({\mathfrak{sl}}_{2})$-modules, see [Dri2]. (As it was shown in [HK] such coboundary structure is related to the category of crystals.)
@@ Acknowledgments
We would like to thank P. Aschieri and D. Bashkirov for stimulating discussions. The authors thank the organisers of the conference “Quantum Groups and Noncommutative Geometry in Prague” (11th to 15th September 2023, Prague, Czech Republic) for warm hospitality and financial support.
@@ 2.2. ${\mathfrak{g}}$-differential spaces and algebras
Take $A={\textstyle{\bigwedge}}{\mathfrak{g}}^{*}$, equipped with the coadjoint action of ${\mathfrak{g}}$ denoted by $L_{x}$ for $x\in{\mathfrak{g}}$. For $x\in{\mathfrak{g}}$ and $f\in{\mathfrak{g}}^{\ast}={\textstyle{\bigwedge}}^{1}{\mathfrak{g}}^{\ast}$ define the contraction operator by $\iota_{x}f=f(x)$. The odd map $i_{x}$ is extended to ${\textstyle{\bigwedge}}{\mathfrak{g}}^{\ast}$ by the super Leibniz rule. Let $e_{a}$ be a basis of ${\mathfrak{g}}$ and $f_{a}$ be the corresponding dual basis in ${\mathfrak{g}}^{\ast}$. The Lie algebra differential on ${\textstyle{\bigwedge}}{\mathfrak{g}}^{\ast}$ may be written as
@@
Suppose that ${\mathfrak{g}}$ has a nondegenerate invariant symmetric bilinear form $B$ (for example, see review in [BKLS]), used to identify ${\mathfrak{g}}\cong{\mathfrak{g}}^{*}$. Let $\operatorname{Cl}({\mathfrak{g}})$ be the Clifford algebra of ${\mathfrak{g}}$ with respect to $B$ define by
@@ 3.1. $\operatorname{\tilde{\sigma}}$-commutators
[…]
@@ 3.2. The $\alpha$ and $\beta$ maps and Lie derivatives
As it was shown in [KP, Lemma 3.7.1] the inner $U_{q}({\mathfrak{sl}}_{2})$-action defined by $\alpha$ coincides with the natural one:
@@
$\beta_{q}(X)=\frac{1+q^{2}}{q}\alpha_{q}(X),\quad\beta_{q}(Y)=\frac{1+q^{2}}{q}\alpha_{q}(Z),\quad\beta_{q}(Z)=\frac{1+q^{2}}{q}\alpha_{q}(Y).$
@@ 3.3. Differential
$\mathrm{d}_{\operatorname{Cl}_{q}}\omega_{q}=[\gamma_{q},\omega]_{\sigma}=\gamma_{q}\omega-(-1)^{p(\omega)}\omega\gamma_{q},\qquad\omega\in\operatorname{Cl}_{q}({\mathfrak{sl}}_{2}).$
@@ 3.4. Contractions
We emphasise that in this case the braided tensor product of algebras in the braided monoidal category of finite-dimensional type 1 $U_{q}({\mathfrak{sl}}_{2})$-modules reduces to the usual tensor product of algebras since the elements of $\operatorname{Cl}(P_{q}({\mathfrak{sl}}_{2}),B_{q})$ are $U_{q}({\mathfrak{sl}}_{2})$-invariant.
@@ 4.3. Quantum exterior algebra
$\mathrm{d}_{\wedge_{q}}=\frac{q^{4}}{1+q^{4}}\left(\frac{1}{c}v_{2}L_{Y}+\frac{q}{(1+q^{2})c}v_{0}L_{Z}+\frac{1}{q^{2}c}v_{-2}L_{X}\right).$

### v2

@@ Abstract.
A. Krutov was supported by the GAČR project 24-10887S and by HORIZON-MSCA-2022-SE-01-01 CaLIGOLA. P. Pandžić was supported by the project “Implementation of cutting-edge research and its application as part of the Scientific Center of Excellence for Quantum and Complex Systems, and Representations of Lie Algebras”, PK.1.1.02, European Union, European Regional Development Fund. This article is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology, www.cost.eu).
@@ 1. Introduction
Let ${\mathfrak{g}}$ be a Lie algebra. H. Cartan introduced the notion of ${\mathfrak{g}}$-differential algebras as a generalisation of algebras of differential forms on manifolds with ${\mathfrak{g}}$-action, [Car1, Car2]. Later ${\mathfrak{g}}$-differential algebras appeared in the study of equivariant cohomology [GS, AM1], in Chern–Weil theory [AM2, Mei], and in relation to (algebraic) Dirac operators and Vogan’s conjecture [AM1, HP1, HP2].
We develop a new approach to quantisation of the notion of ${\mathfrak{g}}$-differential algebras which assumes that the quantum exterior algebra has classical dimension. This is different from previous attempts to generalise the notion of ${\mathfrak{g}}$-differential algebras to the setting of quantum groups and noncommutative geometry; for example, see [AC, SWZ, AS]. Some of these works start with a bicovariant calculus on a quantum group, which usually does not have classical dimension. This is in particular the case for $U_{q}({\mathfrak{sl}}_{2})$, see [Wor], see also [Jur] for the general case. Other works assume the setting of triangular Hopf algebras which is not applicable to the setting of $U_{q}({\mathfrak{sl}}_{2})$ since it is only quasitriangular, see [Dri1].
In this paper we propose a definition of quantised ${\mathfrak{sl}}_{2}$-differential algebras based on the 3-dimensional quantised adjoint $U_{q}({\mathfrak{sl}}_{2})$-module and give first examples, certain quantised Clifford and exterior algebras. The advantage of our approach is that we start with the quantum exterior algebra defined by Berenstein and Zwicknagl [BZ] of the classical dimension instead of a bicovariant calculus. We use the coboundary structure on the category of $U_{q}({\mathfrak{sl}}_{2})$-modules, see [Dri2]. (As it was shown in [HK] such coboundary structure is related to the category of crystals.)
@@ Acknowledgments
We would like to thank P. Aschieri, D. Bashkirov, and R. Ó Buachalla for stimulating discussions. The authors thank the organisers of the conference “Quantum Groups and Noncommutative Geometry in Prague” (11th to 15th September 2023, Prague, Czech Republic) for warm hospitality and financial support. We are also grateful to the referee for useful suggestions.
@@ 2.2. ${\mathfrak{g}}$-differential spaces and algebras
Take $A={\textstyle{\bigwedge}}{\mathfrak{g}}^{*}$, equipped with the coadjoint action of ${\mathfrak{g}}$ denoted by $L_{x}$ for $x\in{\mathfrak{g}}$. For $x\in{\mathfrak{g}}$ and $f\in{\mathfrak{g}}^{\ast}={\textstyle{\bigwedge}}^{1}{\mathfrak{g}}^{\ast}$ define the contraction operator by $\iota_{x}f=f(x)$. The odd map $\iota_{x}$ is extended to ${\textstyle{\bigwedge}}{\mathfrak{g}}^{\ast}$ by the super Leibniz rule. Let $e_{a}$ be a basis of ${\mathfrak{g}}$ and $f_{a}$ be the corresponding dual basis in ${\mathfrak{g}}^{\ast}$. The Lie algebra differential on ${\textstyle{\bigwedge}}{\mathfrak{g}}^{\ast}$ may be written as
@@
Suppose that ${\mathfrak{g}}$ has a nondegenerate invariant symmetric bilinear form $B$ (for example, see review in [BKLS]), used to identify ${\mathfrak{g}}\cong{\mathfrak{g}}^{*}$. Let $\operatorname{Cl}({\mathfrak{g}})$ be the Clifford algebra of ${\mathfrak{g}}$ with respect to $B$ defined by
@@ 3. Motivating example: $\operatorname{Cl}_{q}({\mathfrak{sl}}_{2})$
The quantum exterior algebra ${\textstyle{\bigwedge}}_{q}V_{2\pi}$ and the $q$-deformed Clifford algebra $\operatorname{Cl}_{q}({\mathfrak{sl}}_{2})$ can be considered as results of the deformation quantisation of the classical exterior algebra ${\textstyle{\bigwedge}}{\mathfrak{sl}}_{2}$. The corresponding Poisson structures on ${\textstyle{\bigwedge}}{\mathfrak{sl}}_{2}$ are studied in [AK].
@@ 3.1. $\operatorname{\tilde{\sigma}}$-commutators
$[\omega,\mu]_{\operatorname{\tilde{\sigma}}}=-(-1)^{p(\omega)p(\mu)}[-,-]_{\operatorname{\tilde{\sigma}}}\circ\operatorname{\tilde{\sigma}}(\omega\otimes\mu)\qquad\text{for $\omega,\mu\in\operatorname{Cl}_{q}({\mathfrak{sl}}_{2})$,}$
@@ 3.2. The $\alpha$ and $\beta$ maps and Lie derivatives
As it was shown in [KP, Lemma 3.7.1] the inner $U_{q}({\mathfrak{sl}}_{2})$-action defined by $\alpha_{q}$ coincides with the natural one:
@@
$\beta_{q}(X)=\frac{1+q^{2}}{q}\alpha_{q}(X),\qquad\beta_{q}(Y)=\frac{1+q^{2}}{q}\alpha_{q}(Y),\qquad\beta_{q}(Z)=\frac{1+q^{2}}{q}\alpha_{q}(Z).$
@@
Remark.
One can check that $\beta_{q}$ defines a morphism of “quantum brackets” in the following sense: $[\beta_{q}(x),\beta_{q}(y)]_{\operatorname{\tilde{\sigma}}}=\beta_{q}(\mathop{\mathrm{ad}}\nolimits_{x}y)$ for all $x,y\in{\mathfrak{sl}}_{q}(2)$. Moreover, it follows from Lemma 3.1.1.   Lemma that $\mathop{\mathrm{ad}}\nolimits\circ\operatorname{\tilde{\sigma}}=-\mathop{\mathrm{ad}}\nolimits$ on ${\mathfrak{sl}}_{q}(2)$.
@@ 3.3. Differential
$\mathrm{d}_{\operatorname{Cl}_{q}}\omega_{q}=[\gamma_{q},\omega]_{\operatorname{\tilde{\sigma}}}=\gamma_{q}\omega-(-1)^{p(\omega)}\omega\gamma_{q},\qquad\omega\in\operatorname{Cl}_{q}({\mathfrak{sl}}_{2}).$
@@
Remark.
We note that the differential is dual to the quantised adjoint action of ${\mathfrak{sl}}_{q}(2)$ on itself (a version of “quantum Lie bracket”) in the following sense: $\iota_{x}\iota_{y}\mathrm{d}_{\operatorname{Cl}_{q}}z=\langle\mathop{\mathrm{ad}}\nolimits_{x}y,z\rangle$ for $x,y,z\in{\mathfrak{sl}}_{q}(2)$.
@@ 3.4. Contractions
Remark.
It is subject of further research to describe the derivation property for the action of ${\textstyle{\bigwedge}}_{q}V_{2\pi}$ on $\operatorname{Cl}_{q}({\mathfrak{sl}}_{2})$. We show below that several obvious attempts to ensure this property do not work.
We consider the following possible approach: let us assume that for $x\in{\mathfrak{sl}}_{q}(2)$ and $\omega,\mu\in\operatorname{Cl}_{q}({\mathfrak{sl}}_{2})$,
$\iota_{x}(\omega\mu)=\iota_{x}(\omega)\mu+(-1)^{p(\omega)}\sum_{i}\omega_{i}\iota_{x_{i}}(\mu),$
where $\omega_{i}$ and $x_{i}$ are defined by the following options
  1. (1)
    $\sigma(x\otimes\omega)=\sum\omega_{i}\otimes x_{i}$,
  2. (2)
    $\sigma^{-1}(x\otimes\omega)=\sum\omega_{i}\otimes x_{i}$,
  3. (3)
    $\operatorname{\tilde{\sigma}}(x\otimes\omega)=\sum\omega_{i}\otimes x_{i}$,
In what follows we show that none of these options works in the example of $x=v_{0}$, $\omega=v_{2}$, and $\mu=v_{-2}$. First note that
$\iota_{v_{0}}(v_{2}v_{-2})=\frac{c(1-q^{2})}{q^{2}}v_{0},\qquad\iota_{v_{0}}(v_{2})v_{-2}=0.$
Case 1. We have that
$\sigma(v_{0}\otimes v_{2})=v_{2}\otimes v_{0}+q^{-2}(q^{4}-1)v_{0}\otimes v_{2}.$
Therefore,
$\iota_{x}(\omega)\mu+(-1)^{p(\omega)}\sum_{i}\omega_{i}\iota_{x_{i}}(\mu)=-v_{2}\iota_{v_{0}}(v_{-2})-\frac{q^{4}-1}{q^{2}}v_{0}\iota_{v_{2}}v_{-2}=-\frac{c(q^{4}-1)}{q^{2}}v_{0}\neq\iota_{v_{0}}(v_{2}v_{-2}).$
Case 2. We have that
$\sigma^{-1}(v_{0}\otimes v_{2})=v_{2}\otimes v_{0}.$
Therefore,
$\iota_{x}(\omega)\mu+(-1)^{p(\omega)}\sum_{i}\omega_{i}\iota_{x_{i}}(\mu)=-v_{2}\iota_{v_{0}}(v_{-2})=0\neq\iota_{v_{0}}(v_{2}v_{-2}).$
Case 3. We have that
$\operatorname{\tilde{\sigma}}(v_{0}\otimes v_{2})=\frac{2q^{2}}{1+q^{4}}v_{2}\otimes v_{0}+\frac{q^{4}-1}{1+q^{4}}v_{0}\otimes v_{2}.$
Therefore,
$\iota_{x}(\omega)\mu+(-1)^{p(\omega)}\sum_{i}\omega_{i}\iota_{x_{i}}(\mu)={}$ ${}-\frac{2q^{2}}{1+q^{4}}v_{2}\iota_{v_{0}}(v_{-2})-\frac{q^{4}-1}{1+q^{4}}v_{0}\iota_{v_{2}}v_{-2}$
[…]
Perhaps one could use approach similar to [AC, §6] or consider a braided Hopf algebra structure on ${\textstyle{\bigwedge}}_{q}V_{2\pi}$ similar, for example, to one constructed [KOBS].
In view of the difficulties explained in Remark 3.4.2.   Remark, it is surprising that we have the following theorem.
@@
We emphasise that in this case the braided tensor product of algebras in the braided monoidal category of type 1 finite-dimensional $U_{q}({\mathfrak{sl}}_{2})$-modules reduces to the usual tensor product of algebras since the elements of $\operatorname{Cl}(P_{q}({\mathfrak{sl}}_{2}),B_{q})$ are $U_{q}({\mathfrak{sl}}_{2})$-invariant.
@@ 4. The general definition
It is subject of further research to describe the derivation property for the action of ${\textstyle{\bigwedge}}_{q}V_{2\pi}$ on a quantised ${\mathfrak{sl}}_{2}$-differential algebra; see the discussion in Remark 3.4.2.   Remark.
@@ 4.3. Quantum exterior algebra
In this subsection we show that the quantum exterior algebra of the 3-dimensional quantised adjoint $U_{q}({\mathfrak{sl}}_{2})$-module $V_{2\pi}$ is a quantised ${\mathfrak{sl}}_{2}$-differential algebra in the sense of our definition.
@@
$\mathrm{d}_{\wedge_{q}}=\frac{q^{2}}{1+q^{4}}\left(\frac{1}{c}v_{-2}L_{X}+\frac{q^{3}}{(1+q^{2})c}v_{0}L_{Z}+\frac{q^{2}}{q^{2}}v_{2}L_{Y}\right).$
@@
We note that
$\iota_{v_{2}\wedge v_{0}\wedge v_{-2}}(v_{2}\wedge v_{0}\wedge v_{-2})=\iota_{v_{2}\wedge v_{0}}(cv_{0}\wedge v_{-2})=c^{2}\tfrac{1+q^{2}}{q^{2}}\iota_{v_{2}}v_{-2}=c^{3}\tfrac{1+q^{2}}{q^{2}}.$
Therefore, we can recover the form $B_{q}$ from Remark 3.6.   Remark via contraction operator just as in the classical case of [Kos].